# A Flexible Plug-and-Play Module for Generating Variable-Length

### v1

@@ 5.7.1. Adaptation to Transformer Backbones
Figure 6. The $mAP@K$ evaluation when using ViT as the backbone of CSQ in three datasets. The blue bar denotes CSQ w/o NHL, while yellow denotes CSQ w/ NHL.
5.7.2. Parameter Sensitivity
In Eq.(11), $\lambda$ serves as a hyper-parameter that governs the balance of the long-short cascade self-distillation. We evaluated its values from $\{10^{1},10^{0},10^{-1},10^{-2},10^{-3}\}$ to calculate the $mAP@K$ across three datasets when using CSQ as the deep hashing model. Figure 7 depicts the results and reveals that NHL’s performance exhibits remarkable robustness to variations in $\lambda$. Thus, this finding suggests that the NHL can be seamlessly integrated into many scenarios with minimal adjustment.
@@ 5.7.2. Parameter Sensitivity
Figure 7. The $mAP@K$ results when using various $\lambda$ in three datasets. It shows that NHL is not sensitive to $\lambda$.
(a) CIFAR-10
(b) ImageNet100
(c) MSCOCO
Figure 8. The analysis on more code length settings. The blue bars denote the average ratio of $mAP@K$ at various lengths with and without using NHL. The yellow bars denote the time cost ratio to complete training with and without using NHL
5.7.3. More Code Length Settings
In prior experiments, we configured the hash codes to the five most commonly used lengths within the deep hashing model. This experiment explores the results under a broader range of code length settings. To this end, we established three scenarios for code length:
  • Case 1 set a code length at every 32-bit interval, that is, $b=\{32\times k\}_{k=1}^{4}$ and $m=4$.
  • Case 2 set a code length at every 16-bit interval, that is, $b=\{16\times k\}_{k=1}^{8}$ and $m=8$.
  • Case 3 set a code length at every 8-bit interval, that is, $b=\{8\times k\}_{k=1}^{16}$ and $m=16$.
We use CSQ as the deep hashing model. Figure 8 presents the corresponding results. Here, the blue bars represent the average ratio of $mAP@K$ at various lengths with and without using NHL. The yellow bars indicate the time cost ratio to complete training with and without using NHL. We observe that even with different code length settings, the use of NHL ensures a reduction in overall training time and improves the code quality. Moreover, it is noteworthy that the total training time does not monotonically increase with the number of output code lengths. For instance, the efficiency enhancement ratio in Setting 3 is not as high as in Setting 2. This is attributed to the requirement for the model to undergo more training iterations in Setting 3 to ensure favorable outcomes across a greater number of code lengths.
6. Conclusion
In this paper, we proposed a plug-and-play module NHL for deep hashing models. NHL enables deep hashing models to simultaneously generate hash codes of various lengths, thereby streamlining the training process and reducing the computational burden. Besides, the introduced adaptive weights strategy and the long-short cascade self-distillation ensure the effectiveness of NHL. We conducted extensive experiments on three datasets to evaluate the performance of the NHL. The results demonstrate that NHL accelerates the training process and maintains or enhances retrieval effectiveness across various deep supervised hashing models. In addition to its application in large-scale image retrieval, hashing has also found widespread use in other domains, including cross-modal retrieval (Xie et al., 2020; Tu et al., 2020; Liu et al., 2023), multi-modal retrieval (Zheng et al., 2019; Zhu et al., 2020), and other essential applications (Lu et al., 2023; Zhang et al., 2023; He et al., 2023). Therefore, adapting Nested Hash Layer (NHL) to other domains is a research direction worth exploring.

### v2

@@ 4.4.2 More Code Length Settings
Figure 5: The $mAP@K$ results across three datasets under varying $\lambda$ values demonstrate that NHL is robust to $\lambda$ to some extent.
(d) CIFAR-10
(e) ImageNet100
(f) MSCOCO
Figure 6: Impact of NHL on average mAP@K ratio (blue bars) and training time cost ratio (yellow bars) across various code length settings.
4.5 Compared with Gradient Conflicts Methods and Multi-length Hashing
In this section, we compare our proposed dominance-aware dynamic weighting strategy with other methods aimed at resolving gradient conflicts and two multi-length hashing. Specifically, we evaluate classic gradient conflicts approaches including PCGrad yu2020gradient , GradDrop chen2020just , CAGrad liu2021conflict , and RotoGrad javaloy2022rotograd . The comparison is conducted by replacing the dominance-aware dynamic weighting strategy in our NHL method with the gradient update strategies of these methods and then calculating the $mAP@K$ results across three datasets, using CSQ as the deep hashing model. Besides, we also compared two multi-length hashing models MAH luo2020collaborative and SDMLH nie2022supervised . Table 4 presents the results. It can be observed that, compared to other gradient conflict resolution approaches, our method achieves the best results in more cases. Furthermore, we notice that incorporating these methods often results in negative effects compared to not using any gradient conflict resolution strategy (NHL w/o D). We believe this is due to the nested structure of our parameters, which these classical methods fail to account for adequately. Additionally, the performance of the two multi-code length hashing models, MAH and SDMLH, under most scenarios lags behind that of the CSQ model combined with NHL. This underscores the importance of designing a plug-and-play module, as it allows seamless integration with more advanced deep hashing models to achieve superior retrieval results.
Table 4: The $mAP@K$ results across three datasets using different gradient conflict resolution strategies, with CSQ as the deep semantic hashing model. The best results are highlighted in bold, while the second-best results are underlined.
Method CIFAR-10 ImageNet100 MSCOCO
8 bits 16 bits 32 bits 64 bits 128 bits 8 bits 16 bits 32 bits 64 bits 128 bits 8 bits 16 bits 32 bits 64 bits 128 bits
NHL w/o D 0.707 0.773 0.815 0.823 0.831 0.451 0.803 0.846 0.873 0.884 0.638 0.769 0.822 0.866 0.877
+ PCGrad 0.742 0.781 0.802 0.815 0.830 0.467 0.814 0.852 0.870 0.862 0.632 0.772 0.823 0.852 0.872
+ GradDrop 0.739 0.764 0.805 0.802 0.828 0.461 0.815 0.859 0.872 0.861 0.636 0.776 0.827 0.861 0.875
+ CAGrad 0.720 0.775 0.793 0.811 0.815 0.455 0.815 0.867 0.865 0.873 0.639 0.770 0.825 0.858 0.869
+ RotoGrad 0.735 0.762 0.813 0.819 0.822 0.473 0.808 0.863 0.874 0.877 0.641 0.778 0.830 0.859 0.871
NHL 0.792 0.802 0.818 0.828 0.838 0.495 0.825 0.873 0.880 0.882 0.659 0.778 0.847 0.878 0.881
MAH 0.636 0.649 0.683 0.705 0.714 0.628 0.652 0.689 0.692 0.691 0.567 0.573 0.589 0.599 0.612
SDMLH 0.617 0.684 0.723 0.748 0.763 0.526 0.557 0.618 0.623 0.644 0.602 0.646 0.737 0.761 0.814
5 Conclusion
In this paper, we introduce the plug-and-play module NHL for deep hashing models. NHL allows these models to generate hash codes of different lengths simultaneously, simplifying training and reducing computational load. Additionally, the dominance-aware dynamic weighting strategy and long-short cascade self-distillation enhance NHL’s effectiveness. We performed extensive experiments on three datasets to assess NHL’s performance. The results show that NHL speeds up training while maintaining or improving retrieval effectiveness across various deep supervised hashing models. Our work mainly focuses on common symmetric deep supervised hashing methods, where both the database and query data use the same deep hashing network for generating hash codes. In contrast, NHL is limited to asymmetric deep supervised hashing methods shen2017deep ; jiang2018asymmetric ; chen2019deep ; wu2023deep . They use deep neural networks only for processing query, while direct or indirect optimization of hash codes in the database. Thus, simply changing the hash layer isn’t enough for these methods. We believe that adding optimization designs for the database hash codes could be a promising solution.
Appendix
Appendix A Experimental Settings
A.1 Datasets
We conducted experiments on three widely used datasets in deep hashing for evaluation. CIFAR-10 [20] consists of 60,000 images from 10 classes. Following [3], we randomly select 1,000 images per class as the query set, and 500 images per class as the training set, and use all remaining images as the database. ImageNet100 is a subset of ImageNet [7] with 100 classes. We follow the settings from [9] and randomly select 100 categories. Then, we use all the images of these categories in the training set as the database and the images in the validation set as the queries. Furthermore, we randomly select 13,000 as the training images from the database. MSCOCO [24] is a large-scale multi-label dataset. We consider a subset of 122,218 images from 80 categories, as in previous works [33]. We randomly select 5,000 images from the subset as the query set and use the remaining images as the database. For training, we randomly select 10,000 images from the database. As in most deep hashing settings, two samples are viewed as similar if they correspond to the same label on CIFAR-10 and ImageNet100. For multi-label datasets MSCOCO, two samples are considered similar if they share at least one common label.
A.2 Baselines and Evaluation Metric
We considered the following deep supervised hashing models: DSH [27], DHN [48], DTSH [40], LCDSH [49], DCH [3], DBDH [47], CSQ [45], SHCIR [38], DPN [9], and MDSH [39]. For all the above models, we uniformly adopted ResNet50 [13] as the backbone to extract 2048-dimensional image features.
We employed the mean Average Precision at the top K ($mAP@K$) as the evaluation metric. Specifically, we utilized $mAP@ALL$ for CIFAR-10, $mAP@5000$ for MSCOCO, and $mAP@1000$ for ImageNet100, following the settings used in previous studies [33, 9]. Unless otherwise specified, we set the hash code lengths $b\in\{8,16,32,64,128\}$ for the following experiments, as these lengths are prevalently used in previous works.
A.3 Training Details
For the deep hashing models we adopted, we endeavored to implement all models using PyTorch, based on the code repositories provided in the original papers and the implementation details described therein. The experiments were conducted on a Linux server equipped with 8 NVIDIA GeForce RTX 4090 GPUs. For each model (including those integrated with NHL), a single NVIDIA GeForce RTX 4090 GPU was utilized for both training and testing. Among these models, DHN and DBDH failed to produce valid results on ImageNet100 due to the absence of hyperparameter settings in the original papers, and our grid search method was unable to identify suitable parameters. Additionally, the SHCIR and MDSH models did not propose methods for handling multi-label datasets (e.g., MSCOCO) in their original publications. The batch size $B$ was set to 64. When applying NHL to the deep hashing models, we employed the Adam optimizer [18] and selected the learning rate from $\{10^{-4},10^{-5}\}$. The hyperparameter $\lambda$ was set to $1$. A grid search method was conducted across different scenarios to identify the optimal combination.
Appendix B Proofs
Proposition 1.
[…]
$\alpha_{k}\lVert g_{k}^{(k)}\rVert+\sum_{k<i\leq m}\alpha_{i}cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert\geq 0,$ (10)
then the update of $\mathcal{W}^{(k)}$ is guaranteed to be align-domination.
Proof. Align-domination is satisfied if the inner product between the total gradient $g^{(k)}=\sum_{i=k}^{m}\alpha_{i}g_{i}^{(k)}$ and the dominant gradient $g_{k}^{(k)}$ is non-negative:
$\langle g^{(k)},g_{k}^{(k)}\rangle\geq 0.$
Expanding the inner product:
$\langle g^{(k)},g_{k}^{(k)}\rangle=\alpha_{k}\lVert g_{k}^{(k)}\rVert^{2}+\sum_{k<i\leq m}\alpha_{i}\lVert g_{i}^{(k)}\rVert\lVert g_{k}^{(k)}\rVert\cos\theta_{ik}^{(k)}.$
Dividing through by $\lVert g_{k}^{(k)}\rVert$ (assuming $\lVert g_{k}^{(k)}\rVert>0$), we obtain:
$\alpha_{k}\lVert g_{k}^{(k)}\rVert+\sum_{k<i\leq m}\alpha_{i}\cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert\geq 0.$
Thus, if the inequality holds, the inner product $\langle g^{(k)},g_{k}^{(k)}\rangle\geq 0$, ensuring that the total gradient $g^{(k)}$ is aligned with the dominant gradient $g_{k}^{(k)}$. This confirms align-domination for the update of $\mathcal{W}^{(k)}$.
Proposition 2.
If the following inequality holds for all $k<i\leq m$, then Eq. (10) also holds.
$\alpha_{i}cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert+\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert\geq 0.$ (11)
Proof. For any fixed $k$ and $k<i\leq m$, Eq. (11) can be rewritten as:
$\alpha_{i}\cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert\geq-\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert.$ (12)
This inequality establishes a lower bound on the contribution of each term $\alpha_{i}\cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert$ for $k<i\leq m$. Then, summing Eq. (12) over all $k<i\leq m$, we obtain:
$\sum_{k<i\leq m}\alpha_{i}\cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert\geq\sum_{k<i\leq m}-\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert.$ (13)
Since there are exactly $m-k$ terms in the summation over $k<i\leq m$, the right-hand side simplifies to:
$\sum_{k<i\leq m}-\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert=-(m-k)\cdot\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert=-\alpha_{k}\lVert g_{k}^{(k)}\rVert.$ (14)
Thus, we have:
$\alpha_{k}\lVert g_{k}^{(k)}\rVert+\sum_{k<i\leq m}\alpha_{i}\cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert\geq 0.$ (15)
This proves that Eq. (10) holds whenever Eq. (11) holds for all $k<i\leq m$. ∎
Appendix C The Method to Solve Linear Programming Problem Eq. 5
In the dominance-aware dynamic weighting strategy, we propose the following target for the objective weights $\{\alpha_{k}\}_{k=1}^{m}$:
$\alpha_{i}cos\theta_{ik}^{(k)}\lVert g_{i}^{(k)}\rVert+\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert\geq 0;\ k\leq i\leq m.$ (16)
This section describes how to solve it. Without loss of generality, we first set $\alpha_{1}=1$ for the shortest code’ objective $\mathcal{L}_{1}$, as normalization can subsequently be applied. Then we introduce $\alpha_{i}^{(k)}$ denote only consider to ensure that $L_{i}$ and $L_{k}$ satisfy Eq. (16) on $\mathcal{W}^{(k)}$. Using $cos\theta_{ik}^{(k)}=\frac{g_{i}^{(k)}g_{k}^{(k)}}{\lVert g_{i}^{(k}\rVert\lVert g_{k}^{(k)}\rVert}$ and re-arranging terms, we then get:
$\alpha_{i}^{(k)}(-g_{i}^{(k)}g_{k}^{(k)})\leq\frac{\alpha_{k}}{m-k}\lVert g_{k}^{(k)}\rVert^{2};\ k\leq i\leq m.$ (17)
$\alpha_{i}^{(k)}\leq\frac{\alpha_{k}}{k-m}\frac{\lVert g_{k}^{(k)}\rVert^{2}}{g_{i}^{(k)}g_{k}^{(k)}};\ k\leq i\leq m.$ (18)
$\alpha_{i}^{(k)}=\left\{\begin{array}[]{rcl}1&\text{if}&g_{i}^{(k)}\cdot g_{k}^{(k)}\geq 0\\ \frac{\alpha_{k}}{k-m}\frac{\lVert g_{k}^{(k)}\rVert^{2}}{g_{i}^{(k)}g_{k}^{(k)}}&\text{if}&g_{i}^{(k)}\cdot g_{k}^{(k)}<0;k\leq i\leq m\\ \end{array}\right.$ (19)
$\alpha_{i}=min(\alpha_{i}^{(1)},\alpha_{i}^{(2)},...,\alpha_{i}^{(i)}).$ (20)
It is evident that the computational complexity of calculating $\alpha_{i}^{(k)}$ is $O(lb_{k})$, where $l$ represents the dimension of the data feature $v$. Consequently, the overall computational complexity amounts to $O(lb_{m}m^{2})$, where the $b_{m}$ is the longest code length.
Appendix D The Training Algorithm
In this section, we first present the training algorithm of our proposed NHL in Algorithm 1. Then, we elaborate on the details of our training process. As we discussed in Section 3.5, throughout the training procedure, we monitor the value of each $L_{k}$ and save the model parameters when each $L_{k}$ reaches its minimum to output the corresponding hash codes $h^{(k)}$. In lines 12-14 of Algorithm 1, when a smaller $L_{k}$ is achieved, We record the current model parameters $\theta_{F}$ and $\mathcal{W}$ as the parameters of the model with a length of $b_{k}$, denoted as $\theta_{F}^{(k)}$ and $\mathcal{W}^{(k)}$.
Algorithm 1 The training algorithm with NHL
1:  Initialization: the parameter of the deep hashing model $\{\theta_{F},\mathcal{W}\}$, $\alpha_{k}=1$ for $\forall k$.
2:  repeat
4:     for each $k\in\{1,2,....,m\}$ do
5:        for each $i\in\{1,2,....,m\}$ do
6:           obtain $g_{i}^{(k)}$ by computing standard gradients $g_{i}^{(k)}=\frac{\partial\mathcal{L}_{i}}{\partial\mathcal{W}^{(k)}}$Only calculating the gradients on $\{\mathcal{W}^{(k)}\}_{k=1}^{m}$}
7:        end for
8:     end for
9:     compute $\{\alpha_{k}\}_{k=1}^{m}$ by Eq. (7) and Eq. (6)
10:     renormalize $\{\alpha_{k}\}_{k=1}^{m}$ so that $\sum_{k=1}^{m}\alpha_{k}=m$
12:     if achieved a smaller $\mathcal{L}_{k}$ then
13:        record the current model parameters $\theta_{F}^{(k)}$, $\mathcal{W}^{(k)}$ for the model of length $b_{k}$.
14:     end if
15:  until converged
15:  parameters of deep hashing model $\{\theta_{F}^{(k)}\}_{k=1}^{m}$ and $\{\mathcal{W}^{(k)}\}_{k=1}^{m}$
Appendix E More Experimental Analyses
E.1 Hyper-parameters
We further conducted experimental validation on the CSQ model under different learning rate settings. We set the learning rates to $\{10^{-1},10^{-2},10^{-3},10^{-4},10^{-5},10^{-6}\}$ and performed experiments across three datasets. Figure 7 presents the results. Based on this experiment, we determined the optimal learning rate range to be $\{10^{-4},10^{-5}\}$.
@@ E.1 Hyper-parameters
Figure 7: The mAP@K results across three datasets under different learning rates.
E.2 Backbone Analysis
The NHL is designed for versatile integration with various deep supervised hashing models. To rigorously validate the generalizability of NHL, we extended this investigation by incorporating it into the CSQ deep hashing model while utilizing a diverse range of alternative backbone architectures for feature extraction. Specifically, we experimented with ResNet variants of different scales, including ResNet18 and ResNet34. We also present the results of ResNet50 used previously for comparison. Furthermore, to assess NHL’s compatibility with distinct network structures, we also employed the Vision Transformer (ViT_B_16) [8] and the lightweight MobileNetV2 architecture [35]. As demonstrated in Table 5, NHL consistently maintains its effectiveness when paired with these varied backbones. This underscores NHL’s broad applicability and its robustness in enhancing hashing performance across different types and scales of feature extractors, from various CNNs to transformer-based models.
Table 5: The $mAP@K$ evaluation when using different feature extraction networks as the backbone of CSQ across three datasets. We compared the results of using NHL (w/ NHL) with those of not using NHL (w/o NHL) on these three datasets.
Data Backbone w/o NHL (The Original Model) w/ NHL
CIFAR-10 (mAP@ALL) ResNet18 0.654 0.710 0.743 0.717 0.739 0.712 0.664 0.718 0.762 0.793 0.817 0.750 $(+5.36\%)$
ResNet34 0.682 0.700 0.724 0.730 0.722 0.711 0.691 0.724 0.757 0.787 0.826 0.757 $(+6.37\%)$
ResNet50 0.762 0.786 0.798 0.798 0.807 0.790 0.792 0.802 0.818 0.828 0.838 0.816 $(+3.21\%)$
MobileNetV2 0.661 0.725 0.750 0.774 0.788 0.739 0.677 0.735 0.779 0.799 0.816 0.761 $(+2.92\%)$
ViT_B_16 0.891 0.903 0.907 0.912 0.913 0.905 0.895 0.909 0.919 0.931 0.937 0.918 $(+1.44\%)$
ImageNet100 (mAP@1000) ResNet18 0.375 0.727 0.754 0.773 0.795 0.684 0.413 0.722 0.774 0.804 0.815 0.705 $(+3.03\%)$
ResNet34 0.398 0.773 0.815 0.836 0.839 0.732 0.465 0.784 0.828 0.850 0.855 0.756 $(+3.31\%)$
ResNet50 0.456 0.822 0.860 0.877 0.878 0.778 0.495 0.825 0.873 0.880 0.882 0.787 $(+1.59\%)$
MobileNetV2 0.414 0.679 0.745 0.788 0.802 0.685 0.422 0.708 0.759 0.800 0.819 0.701 $(+2.33\%)$
ViT_B_16 0.523 0.888 0.910 0.915 0.915 0.830 0.533 0.887 0.906 0.918 0.925 0.834 $(+0.43\%)$
MSCOCO (mAP@5000) ResNet18 0.532 0.653 0.745 0.792 0.811 0.706 0.576 0.679 0.760 0.807 0.824 0.729 $(+3.19\%)$
ResNet34 0.554 0.715 0.789 0.830 0.836 0.744 0.591 0.731 0.802 0.840 0.847 0.762 $(+2.33\%)$
ResNet50 0.596 0.750 0.847 0.877 0.871 0.788 0.659 0.778 0.847 0.878 0.881 0.809 $(+2.58\%)$
MobileNetV2 0.543 0.668 0.742 0.806 0.823 0.716 0.578 0.691 0.766 0.815 0.836 0.737 $(+2.90\%)$
ViT_B_16 0.638 0.790 0.883 0.895 0.906 0.822 0.675 0.794 0.889 0.898 0.902 0.831 $(+1.12\%)$
Appendix F Ablation Study
This section presents the comprehensive results of NHL variants applied to deep hashing models. Figures 8,  6, and  7 showcase the outcomes for CIFAR-10, ImageNet100, and MSCOCO, respectively. The best results are highlighted in bold, while the second-best results are underlined. From these results, it is evident that in the vast majority of cases, employing the complete NHL method achieves optimal performance. Additionally, in certain scenarios, using variants of NHL yields the best results, which can be attributed to the inherent differences among various deep hashing models, as well as the influence of different hash code lengths and datasets. As a plug-and-play module, NHL demonstrates sufficient robustness and adaptability across diverse applications.
Table 6: The mAP@K comparison results on the ImageNet100 dataset when different deep hashing models use the original hash layer, NHL, or the variants of NHL. The best results are highlighted in bold, while the second-best results are underlined.
Data Length NHL DSH DTSH LCDSH DCH CSQ DPN SHCIR MDSH
ImageNet100 (mAP@1000) 8bit w/o 0.703 0.432 0.248 0.776 0.456 0.436 0.789 0.785
-basic 0.630 0.544 0.407 0.809 0.424 0.453 0.744 0.697
w/o D 0.688 0.512 0.413 0.810 0.451 0.456 0.764 0.787
w/o L 0.663 0.535 0.426 0.797 0.487 0.453 0.777 0.756
w/ 0.755 0.552 0.422 0.809 0.495 0.487 0.798 0.794
16bit w/o 0.808 0.710 0.395 0.834 0.822 0.827 0.861 0.845
-basic 0.787 0.682 0.571 0.845 0.781 0.830 0.831 0.812
w/o D 0.797 0.701 0.597 0.846 0.803 0.832 0.822 0.836
w/o L 0.784 0.721 0.554 0.839 0.820 0.825 0.828 0.827
w/ 0.816 0.714 0.568 0.842 0.825 0.829 0.881 0.851
32bit w/o 0.827 0.770 0.542 0.845 0.860 0.864 0.879 0.874
-basic 0.814 0.765 0.622 0.856 0.847 0.862 0.867 0.859
w/o D 0.808 0.721 0.645 0.858 0.846 0.861 0.867 0.846
w/o L 0.807 0.749 0.601 0.850 0.848 0.858 0.861 0.860
w/ 0.829 0.766 0.628 0.855 0.873 0.860 0.889 0.878
64bit w/o 0.828 0.784 0.608 0.859 0.877 0.870 0.883 0.895
-basic 0.825 0.773 0.671 0.864 0.874 0.875 0.881 0.879
w/o D 0.818 0.760 0.679 0.863 0.873 0.874 0.885 0.872
w/o L 0.820 0.769 0.631 0.854 0.864 0.870 0.882 0.879
w/ 0.838 0.788 0.657 0.860 0.880 0.877 0.893 0.884
128bit w/o 0.822 0.803 0.692 0.848 0.878 0.877 0.881 0.894
-basic 0.827 0.782 0.703 0.863 0.888 0.878 0.887 0.888
w/o D 0.820 0.764 0.714 0.866 0.884 0.880 0.884 0.883
w/o L 0.825 0.770 0.665 0.857 0.878 0.877 0.887 0.886
w/ 0.841 0.792 0.692 0.863 0.882 0.881 0.898 0.896
Table 7: The $mAP@K$ comparison results on the MSCOCO dataset when different deep hashing models use the original hash layer, NHL, or the variants of NHL. The best results are highlighted in bold, while the second-best results are underlined.
Data Length NHL DSH DHN DTSH LCDSH DCH DBDH CSQ DPN
MSCOCO (mAP@1000) 8bit w/o 0.685 0.659 0.706 0.687 0.695 0.655 0.596 0.575
-basic 0.693 0.714 0.737 0.708 0.720 0.676 0.635 0.621
w/o D 0.662 0.705 0.738 0.727 0.721 0.690 0.638 0.633
w/o L 0.688 0.708 0.734 0.703 0.692 0.688 0.636 0.646
w/ 0.714 0.724 0.751 0.713 0.723 0.692 0.659 0.638
16bit w/o 0.722 0.751 0.770 0.769 0.756 0.727 0.750 0.757
-basic 0.730 0.757 0.769 0.757 0.768 0.732 0.735 0.747
w/o D 0.705 0.751 0.775 0.771 0.773 0.737 0.769 0.779
w/o L 0.726 0.764 0.789 0.761 0.736 0.747 0.749 0.774
w/ 0.735 0.760 0.793 0.773 0.769 0.748 0.778 0.769
32bit w/o 0.757 0.786 0.810 0.787 0.762 0.760 0.847 0.828
-basic 0.747 0.789 0.791 0.774 0.788 0.764 0.822 0.820
w/o D 0.731 0.783 0.805 0.788 0.787 0.765 0.822 0.835
w/o L 0.749 0.792 0.809 0.789 0.757 0.781 0.811 0.819
w/ 0.764 0.794 0.819 0.794 0.786 0.778 0.847 0.837
64bit w/o 0.779 0.810 0.823 0.825 0.777 0.769 0.877 0.862
-basic 0.765 0.809 0.796 0.780 0.791 0.783 0.866 0.853
w/o D 0.749 0.804 0.808 0.798 0.789 0.783 0.866 0.860
w/o L 0.765 0.815 0.814 0.826 0.765 0.803 0.857 0.849
w/ 0.789 0.837 0.823 0.828 0.789 0.809 0.881 0.872
128bit w/o 0.769 0.832 0.831 0.836 0.734 0.800 0.871 0.863
-basic 0.773 0.820 0.797 0.786 0.787 0.799 0.876 0.868
w/o D 0.761 0.822 0.814 0.802 0.787 0.797 0.877 0.872
w/o L 0.774 0.828 0.815 0.839 0.764 0.811 0.873 0.866
w/ 0.789 0.837 0.823 0.828 0.789 0.809 0.881 0.872
Table 8: The mAP@K comparison results on the CIFAR-10 dataset when different deep hashing models use the original hash layer, NHL, or the variants of NHL. The best results are highlighted in bold, while the second-best results are underlined.
Data Length NHL DSH DHN DTSH LCDSH DCH DBDH CSQ DPN SHCIR MDSH
CIFAR-10 (mAP@1000) 8bit w/o 0.690 0.718 0.754 0.715 0.776 0.737 0.762 0.703 0.754 0.755
-basic 0.598 0.724 0.560 0.759 0.779 0.731 0.652 0.731 0.761 0.752
w/o D 0.620 0.730 0.772 0.751 0.772 0.750 0.707 0.727 0.740 0.743
w/o L 0.641 0.752 0.747 0.773 0.781 0.737 0.769 0.632 0.760 0.760
w/ 0.717 0.739 0.766 0.775 0.787 0.748 0.792 0.729 0.758 0.762
16bit w/o 0.731 0.765 0.778 0.771 0.802 0.771 0.786 0.757 0.791 0.808
-basic 0.688 0.778 0.702 0.782 0.806 0.769 0.758 0.768 0.796 0.788
w/o D 0.667 0.769 0.785 0.766 0.798 0.770 0.773 0.783 0.777 0.775
w/o L 0.710 0.797 0.775 0.785 0.808 0.764 0.777 0.752 0.793 0.801
w/ 0.732 0.776 0.790 0.799 0.810 0.785 0.802 0.765 0.797 0.811
32bit w/o 0.740 0.813 0.799 0.817 0.829 0.796 0.798 0.790 0.820 0.829
-basic 0.726 0.818 0.777 0.815 0.827 0.797 0.795 0.804 0.834 0.823
w/o D 0.733 0.783 0.814 0.788 0.822 0.799 0.815 0.812 0.812 0.814
w/o L 0.739 0.815 0.796 0.799 0.828 0.793 0.803 0.792 0.828 0.831
w/ 0.744 0.824 0.802 0.825 0.833 0.804 0.818 0.795 0.824 0.838
64bit w/o 0.727 0.837 0.831 0.826 0.830 0.829 0.798 0.804 0.844 0.844
-basic 0.734 0.834 0.810 0.823 0.834 0.822 0.812 0.819 0.853 0.845
w/o D 0.740 0.802 0.820 0.803 0.827 0.811 0.823 0.817 0.834 0.838
w/o L 0.739 0.833 0.808 0.814 0.835 0.816 0.821 0.819 0.848 0.849
w/ 0.743 0.835 0.822 0.839 0.843 0.825 0.828 0.826 0.849 0.852
128bit w/o 0.381 0.853 0.811 0.854 0.825 0.829 0.807 0.819 0.828 0.832
-basic 0.752 0.846 0.835 0.835 0.829 0.838 0.824 0.824 0.827 0.852
w/o D 0.744 0.821 0.831 0.824 0.831 0.834 0.831 0.821 0.838 0.845
w/o L 0.736 0.852 0.818 0.828 0.839 0.873 0.832 0.827 0.857 0.858
w/ 0.749 0.866 0.836 0.868 0.844 0.844 0.838 0.824 0.848 0.861